20220617 / restart in progress

EDIT 14h00:

  • ici se trouve une bordée d'injure que la bienséance nous interdit de reproduire
  • we will have to redo it all over again next week… 'yeah… \o/'
  • data13 (/Xnfs/cosmos, /Xnfs/zoomin) : IN PROGRESS, delivered hardware has issues

EDIT 11h30:

  • reboot mostly done
  • some deb9/sge queues still powered off (see qstat)
  • some Lake nodes still powered off (network & hardware modifications TODO)
  • all deb11/slurm active
  • data13 (/Xnfs/cosmos, /Xnfs/zoomin) : IN PROGRESS, add hardware and reboot

At that time, mostly upgrades on various servers. Expect deconnections.

  • data10 (chimie) : done (upgrades on storage and kernel)
  • ssh.psmn : done (no reboot needed)
  • allo-psmn : done (reboot needed)
newsfeed/20220617.txt · Dernière modification: 2022/06/17 15:10 par ltaulell